OpenStudy (anonymous):

Solve |3x - 4| = 6 i got {-10/3, 10/3}?

OpenStudy (anonymous):

how do u solve?

OpenStudy (anonymous):

x=-2/3,10/3

OpenStudy (anonymous):

the key is that |x|=a x=+/- a

OpenStudy (anonymous):

I'm afraid your answer is wrong..... But don't worry I will solve your Confusions. :)

OpenStudy (anonymous):

\[|3x−4|=6\] \[\implies \]\[3x−4=6\] or \[3x−4=−6\] So you have to solve both of those two equations to get your solutions.

OpenStudy (anonymous):

You solved the first one correctly to get \(x = \frac{10}{3}\) Now solve the second one to find the other solution.

OpenStudy (anonymous):

|3x−4|=6 So, 3x−4=6 or 3x−4=−6. 1st 3x=10 --> x=10/3 or 2nd 3x=-2 --> x=-2/3